# ACDC: The Adverse Conditions Dataset with Correspondences for Robust Semantic Driving Scene Perception

### v4

@@ Abstract
Most of the prominent large-scale image-based datasets for driving scene perception, including Cityscapes [1], Vistas [2] and KITTI [3], are dominated by images captured under normal visual conditions, i.e., at daytime and in clear weather. Yet, vision applications such as automated driving impose a strict requirement on perception algorithms to maintain satisfactory performance in adverse domains. Although there have been efforts to include adverse visual domains in large-scale datasets, such as Oxford RobotCar [4] and BDD100K [5], these efforts focus either on localization/mapping tasks [4, 6] or on recognition tasks which do not involve dense pixel-level outputs, such as object detection [5, 7, 8]. For instance, while a notable 40% of the object detection set of BDD100K pertains to nighttime, only 3% of the images in its 10K semantic segmentation set, namely 345 images, are captured at nighttime [9]. In addition, the pixel-level annotation process for adverse-condition images is kept identical in [5, 10] to the normal-condition case, which leads to errors in the ground truth and renders it unreliable [9]. In contrast, seminal previous work [1] has underlined the need for specialized techniques and datasets for pixel-level semantic perception in adverse visual conditions, due to the inherent aleatory uncertainty in images captured in such conditions. These render entire image regions indiscernible even for humans.
@@
ACDC constitutes a response to this need for a large-scale driving dataset specialized to adverse conditions, in terms of (i) size, (ii) domain adversity, and (iii) featured tasks. ACDC includes 5509 images with high-quality pixel-level panoptic annotations. From this complete set of images, 4006 images are distributed equally among four common adverse conditions in real-world driving environments, namely fog, nighttime, rain, and snow, and the rest 1503 pertain to normal conditions, i.e. daytime and clear weather, thus granting ACDC a scale that slightly exceeds that of Cityscapes. The adverse-condition part of the dataset was deliberately recorded with the respective adverse conditions clearly present. Thus, a large domain shift from the normal clear-weather daytime conditions was achieved. Moreover, for each adverse-condition image, a corresponding normal-condition image of the same scene from approximately the same viewpoint is provided, intended for use by weakly supervised methods.
@@
As to the tasks that our dataset supports, apart from standard dense semantic perception tasks such semantic segmentation, object detection, instance segmentation and panoptic segmentation, we add the task of uncertainty-aware semantic segmentation. For the latter we introduce a specialized annotation protocol and a dedicated performance metric, termed average uncertainty-aware IoU (AUIoU). The key characteristic of uncertainty-aware semantic segmentation is the principled inclusion of image regions with indiscernible semantic content—invalid regions—in annotation and evaluation. In particular, the annotation protocol for our adverse-condition images leverages privileged information in the form of the corresponding normal-condition images and the original adverse-condition videos, which enables to reliably assign legitimate semantic labels to invalid regions and to include them in the evaluation both for the aforementioned standard semantic perception tasks and for uncertainty-aware semantic segmentation. For the latter task, the separation of labeled pixels into invalid and valid is encoded in a binary mask. While both tasks require a hard semantic prediction, the uncertainty-aware task additionally expects a confidence map prediction. AUIoU is designed to take into account both the semantic and the confidence prediction and to reward predictions with low confidence on invalid pixels and high confidence on valid pixels. The requirement for an additional confidence prediction is relevant for safety-oriented applications, as it can help the downstream decision-making system avoid the fatal consequences of a low-confidence prediction being false, e.g. when a pedestrian is missed.
@@
Overall, we experiment with ACDC on all five semantic perception tasks it supports in four main directions: (i) unsupervised and weakly supervised normal-to-adverse and sensor-level domain adaptation, (ii) supervised semantic perception in adverse conditions, (iii) evaluation of models externally pre-trained on normal conditions, and (iv) evaluation of uncertainty-aware semantic segmentation baselines and oracles. Results show that access to ground-truth annotations under adverse conditions is indispensable for achieving high performance, as pre-trained models severely deteriorate under adverse conditions. Moreover, the real-world CityscapesACDC adaptation scenario stands for a challenging setting for all state-of-the-art UDA methods, which still trail fully supervised counterparts. This underlines the need for UDA methods that perform better when handling adverse target domains and highlights the importance of ACDC in steering future work in this direction. Finally, the uncertainty-aware annotations of ACDC create significant room for improvement over simple confidence prediction baselines and help promote future work on semantic segmentation methods that simultaneously models uncertainty.
@@ 2 Related Work
Datasets for driving scene perception include real-world and synthetic sets that support geometric and recognition tasks. KITTI [3] and Cityscapes [1] pioneered this area with LiDAR and semantic image annotations, respectively. Subsequent datasets mostly aimed at increasing the scale [49], diversity [2] and number of tasks [5]. As high-quality pixel-level annotations proved hard to acquire [1, 2], another line of work focused on creating synthetic sets at an even larger scale [35, 34, 50, 51, 52] and in which ground truth is automatically generated, as well as translating real datasets to adverse conditions such as fog or rain [36, 38, 53]. Oxford Robotcar [4] was the first real-world large-scale dataset in which adverse visual conditions such as nighttime, rain and snow were significantly represented, but it did not feature semantic annotations. While more recent large-scale sets [54, 55] that cover adverse conditions, such as Waymo Open [7] and nuScenes [8], include bounding boxes, they still lack dense pixel-level semantic annotations, which are vital for real-world autonomous agents [56]. BDD100K [5] was the only exception to this rule prior to the publication of the conference version [37] of this paper, with ca. 13% of its 10000 pixel-level annotations pertaining to adverse conditions but containing severe errors [9]. At the same time, only a small portion of each of the 1881 adverse-condition images in ADUULM [57] is annotated. On the other hand, several sets with small-scale pixel-level annotations covering adverse conditions [10] had been presented, focusing on fog [36, 58], nighttime [39, 9], and rain [59]. A notable case is Dark Zurich [9], with 201 fine pixel-level nighttime annotations and a dedicated annotation protocol and evaluation metric that handles regions with ambiguous content. The initial, conference version of ACDC improved both upon BDD100K, in terms of ground truth quality, and Dark Zurich, in terms of scale and condition diversity, featuring 5509 high-quality fine instance-level semantic annotations in which fog, night, rain, snow, and normal conditions are evenly represented. Since the initial publication of ACDC, a larger-scale version of WildDash, namely Wilddash2 [60], has been released. The present extended version of ACDC exceeds the scale of Wilddash2 annotations (5509 vs. 5000). Moreover, in contrast both to Wilddash2 and to other recent dense semantic perception datasets with adverse conditions [61, 62], ACDC features cross-condition image-level correspondences with normal-condition reference images as well as a specialized annotation protocol which hinges exactly on the aforementioned correspondences to allow the assignment of legitimate, reliable semantic labels to indiscernible image regions which would otherwise be impossible to label.
@@
Instance segmentation and panoptic segmentation distinguish instances in the images compared to the semantic segmentation task. Instance segmentation aims to assign a segmentation mask for each object of interest in the image, while panoptic segmentation task encompasses masks for both stuff and things classes as the combination of semantic segmentation and instance segmentation. Both of these two tasks are applied in various real-world applications like robotics, healthcare and geoscience. For instance segmentation, many works either follow the detect-then-segment pipeline [86] or explore to generate a dynamic number of instance masks directly with clustering or dynamic kernels. The pioneering work, Mask R-CNN generate instance masks based on object bounding box proposals. The following specialized instance segmentation works like Cascaded Mask R-CNN, HTC and Detectors mainly work on extracting better feature representations and generating more accurate object proposals. For panoptic segmentation, early works strive to combine semantic segmentation models and instance segmentation models to predict unified panoptic masks. Recent works begin to view the panoptic segmentation task in a unified perspective and formulate the stuff and things segmentation as a set prediction problem. The current state of the art for panoptic segmentation includes i.a. PanopticFPN and Panoptic-Deeplab with specialized instance prediction branches and semantic prediction branches, K-Net with dynamic kernels, and MaskFormer and Mask2Former with transformer architecture. Although current instance segmentation and panoptic segmentation models obtains impressive improvement on the popular Cityscapes benchmark, we present that these state-of-the-art generalize poorly to ACDC. The adverse images in ACDC poses a new challenges for instance segmentation models and panoptic segmentation models and would be able to further encourage the development in the fields.
@@
Adaptation of object detection networks from a labeled source domain to another unlabeled target domain is also an active research field. Currently, there are mainly two categories of UDA for object detection methods: domain alignment and self-training. Domain alignment strive to bridge the domain gap by minimizing the domain discrepancy through style transfer, adversarial training and graph matching. Self-training relies on pseudo labels to extract rich knowledge contained in target domain and presents promising performance. Although all aforementioned have been proven effective in alleviating the damage of domain shifts, due to the lack of dataset with rich adverse conditions, these methods are mainly evaluated on synthetic-to-real, cross-city and cross-camera settings. The normal-to-adverse adaptation, which is of high value for intelligent driving systems, is rarely discussed. ACDC provides a large-scale data in adverse conditions and would promote the progress of unsupervised and weakly supervised object detection methods for adverse conditions.
@@ 3 ACDC Dataset
We base the design of ACDC on the same general principles as seminal normal-condition datasets [1] and adapt the collection and annotation process to fit better the adverse condition setting at hand.
@@ 3.1 Collection
Thus, we recorded several days of video in Switzerland by driving around in a car, primarily in urban areas but also on highways and in rural regions. In order to have a clear domain separation between different adverse conditions, we use the following criterion for the adverse-condition recordings: each recording takes place under only one type of adversity from a set of four items, i.e., fog, nighttime, rain, and snow. For example, our foggy recordings are performed at daytime and without rain or snow. For snow, both snowfall and snow cover on the ground are admissible. Moreover, we keep for further processing only the parts of the adverse-condition recordings that correspond to an intense presence of the respective condition, so as to maximize the domain shift from normal conditions as well as domain adversity.
@@ 3.2 Correspondence Establishment
Our camera also provides GPS readings, which allow us to establish image-level correspondences between adverse-condition and normal-condition recordings. In particular, for each adverse-condition recording, we perform a normal-condition recording along exactly the same route. We then use the sequences of GPS measurements of the two recordings to perform a global dynamic-programming-based matching of the adverse GPS sequence to the normal one, where the objective is defined by the Euclidean distances of matched pairs of GPS samples. Our global matching handles routes with loops better than simple nearest neighbors. Each adverse-condition frame is then matched to a normal-condition frame based on the corresponding matched samples of the GPS sequences.
@@ 3.4 Semantic Annotation
Our annotation protocol for the 4006 adverse-condition images consists of two cascaded annotation stages. At stage 1, a semantic labeling draft is manually produced from the adverse-condition image , in which pixels that cannot be unquestionably assigned to a single semantic class are left unlabeled. At stage 2, the corresponding normal-condition image and the adverse-condition video from which was extracted are used to augment and finalize the annotation. In particular, the annotator can assign a legitimate label to pixels that were left unlabeled in stage 1 and correct pixels that were incorrectly labeled in stage 1. Pixels that remain unclear in stage 2 are left unlabeled and are not used for training or evaluation.
@@
The 4006 fine pixel-level annotations of ACDC were created by a professional team of annotators to ensure high-quality ground truth. Annotators were asked to be conservative in labeling pixels in both stages, so as to minimize errors. Both the initial draft from stage 1 and the final annotation from stage 2 passed through quality control. The total time required for annotating a single adverse-condition image was 3.3 h on average. The semantic annotation of the 1503 normal-condition images is conducted in the standard way by using only the input normal-condition image.
The class specifications of ACDC are directly inherited from Cityscapes. In particular, we annotate the 19 evaluation classes of Cityscapes, which include the most common and traffic-related objects in driving scenes. Objects that belong to classes outside this set receive a fall-back label and are not used for training or evaluation. This choice of classes provides full compatibility of ACDC to Cityscapes and other normal-condition datasets for semantic segmentation [2, 5]. Detailed annotation statistics are presented in Fig. 1. An example of our two-stage annotation protocol is shown in Fig. 2 for a snowy image. Note the assignment of a region in the lower right part of the image that is unlabeled at stage 1 (Fig. 2b) to the road label at stage 2 (Fig. 2d), thanks to the clear view from the normal-condition image.

### v5

@@ Abstract
Most of the prominent large-scale image-based datasets for driving scene perception, including Cityscapes [1], Vistas [2] and KITTI [3], are dominated by images captured under normal visual conditions, i.e., at daytime and in clear weather. Yet, vision applications such as automated driving impose a strict requirement on perception algorithms to maintain satisfactory performance in adverse domains. Although there have been efforts to include adverse visual domains in large-scale datasets, such as Oxford RobotCar [4] and BDD100K [5], these efforts focus either on localization/mapping tasks [4, 6] or on recognition tasks which do not involve dense pixel-level outputs, such as object detection [5, 7, 8]. For instance, while a notable 40% of the object detection set of BDD100K pertains to nighttime, only 3% of the images in its 10K semantic segmentation set, namely 345 images, are captured at nighttime [9]. Thus, at the time when the conference version of this article appeared, there was no large adverse-condition pixel-level segmentation dataset available. In addition, the pixel-level annotation process for adverse-condition images is kept identical in [5, 10] to the normal-condition case, which leads to errors in the ground truth and renders it unreliable [9]. The application of highly automated annotation pipelines with little human intervention, which has recently catalyzed the creation of large-scale pixel-level segmentation data [11], to adverse-condition images is thus especially prone to creating severe annotation errors, which is detrimental for model training, let alone rigorous evaluation, which requires highly accurate ground truth as reference. In contrast, seminal previous work [1] has underlined the need for specialized techniques and datasets for pixel-level semantic perception in adverse visual conditions, due to the inherent aleatory uncertainty in images captured in such conditions. These render entire image regions indiscernible even for humans.
@@
ACDC constitutes precisely a response to this need for a large-scale labeled driving segmentation dataset specialized to adverse conditions, in terms of (i) size, (ii) domain adversity, and (iii) featured tasks. ACDC includes 5509 images with high-quality pixel-level panoptic annotations. From this complete set of images, 4006 images are distributed equally among four common adverse conditions in real-world driving environments, namely fog, nighttime, rain, and snow, and the rest 1503 pertain to normal conditions, i.e. daytime and clear weather, thus granting ACDC a scale that slightly exceeds that of Cityscapes. The adverse-condition part of the dataset was deliberately recorded with the respective adverse conditions clearly present. Thus, a large domain shift from the normal clear-weather daytime conditions was achieved. Moreover, for each adverse-condition image, a corresponding normal-condition image of the same scene from approximately the same viewpoint is provided, intended for use by weakly supervised methods.
@@
As to the tasks that our dataset supports, apart from standard dense semantic perception tasks such as semantic segmentation, object detection, instance segmentation and panoptic segmentation, we add the task of uncertainty-aware semantic segmentation. For the latter we introduce a specialized annotation protocol and a dedicated performance metric, termed average uncertainty-aware IoU (AUIoU). The key characteristic of uncertainty-aware semantic segmentation is the principled inclusion of image regions with indiscernible semantic content—invalid regions—in annotation and evaluation. In particular, the annotation protocol for our adverse-condition images leverages privileged information in the form of the corresponding normal-condition images and the original adverse-condition videos, which enables us to reliably assign legitimate semantic labels to invalid regions and to include them in the evaluation both for the aforementioned standard semantic perception tasks and for uncertainty-aware semantic segmentation. For the latter task, the separation of labeled pixels into invalid and valid is encoded in a binary mask. While both tasks require a hard semantic prediction, the uncertainty-aware task additionally expects a confidence map prediction. AUIoU is designed to take into account both the semantic and the confidence prediction and to reward predictions with low confidence on invalid pixels and high confidence on valid pixels. The requirement for an additional confidence prediction is relevant for safety-oriented applications, as it can help the downstream decision-making system avoid the fatal consequences of a low-confidence prediction being false, e.g. when a pedestrian is missed.
Compared to previous related datasets, ACDC thus boasts: (i) several different conditions, all evenly represented, instead of a single condition or a highly skewed distribution over conditions, (ii) a high representation of adversities in the set instead of a bias towards normal conditions, (iii) large-scale adverse-condition data instead of small-scale ones, (iv) reliable high-quality large-scale annotations in adverse conditions instead of unreliable ones [5] in this setting, (v) legitimate labels even for invalid regions which are indiscernible even for humans thanks to privileged correspondence-based information instead of mere exclusion of these regions from the ground truth. A quantitative comparison to other adverse-condition datasets across these axes is deferred to Sec. 3.6.
@@
Overall, we experiment with ACDC on all five semantic perception tasks it supports in four main directions: (i) unsupervised and weakly supervised normal-to-adverse and sensor-level domain adaptation, (ii) supervised semantic perception in adverse conditions, (iii) evaluation of models externally pre-trained on normal conditions, and (iv) evaluation of uncertainty-aware semantic segmentation baselines and oracles. Results show that access to ground-truth annotations under adverse conditions is indispensable for achieving high performance, as pre-trained models severely deteriorate under adverse conditions. Moreover, the real-world CityscapesACDC adaptation scenario stands for a challenging setting for all state-of-the-art UDA methods, which still trail fully supervised counterparts. This underlines the need for UDA methods that perform better when handling adverse target domains and highlights the importance of ACDC in steering future work in this direction. Finally, the uncertainty-aware annotations of ACDC create significant room for improvement over simple confidence prediction baselines and help promote future work on semantic segmentation that simultaneously models uncertainty.
@@ 2 Related Work
Datasets for driving scene perception include real-world and synthetic sets that support geometric and recognition tasks. KITTI [3] and Cityscapes [1] pioneered this area with LiDAR and semantic image annotations, respectively. Subsequent datasets mostly aimed at increasing the scale [50], diversity [2] and number of tasks [5]. As high-quality pixel-level annotations proved hard to acquire [1, 2], another line of work focused on creating synthetic sets at an even larger scale [36, 35, 51, 52, 53] and in which ground truth is automatically generated, as well as translating real datasets to adverse conditions such as fog or rain [37, 39, 54]. Oxford Robotcar [4] was the first real-world large-scale dataset in which adverse visual conditions such as nighttime, rain and snow were significantly represented, but it did not feature semantic annotations. While more recent large-scale sets [55, 56] that cover adverse conditions, such as Waymo Open [7] and nuScenes [8], include bounding boxes, they still lack dense pixel-level semantic annotations, which are vital for real-world autonomous agents [57] and are featured by ACDC. BDD100K [5] was the only exception to this rule prior to the publication of the conference version [38] of this paper, with ca. 13% of its 10000 pixel-level annotations pertaining to adverse conditions but containing severe errors [9], contrary to the high-quality reliable ground truth of ACDC. At the same time, only a small portion of each of the 1881 adverse-condition images in ADUULM [58] is annotated, whereas 92.5% of the pixels in ACDC are annotated. On the other hand, several sets with small-scale pixel-level annotations covering adverse conditions [10] had been presented, focusing on fog [37, 59], nighttime [40, 9], and rain [60]. A notable case is Dark Zurich [9], with 201 fine pixel-level nighttime annotations and a dedicated annotation protocol and evaluation metric that handles regions with ambiguous content. The initial, conference version of ACDC improved both upon BDD100K, in terms of ground truth quality, and Dark Zurich, in terms of scale and condition diversity, featuring 5509 high-quality fine instance-level semantic annotations in which fog, night, rain, snow, and normal conditions are evenly represented. Since the initial publication of ACDC, a larger-scale version of WildDash, namely Wilddash2 [61], has been released. The present extended version of ACDC exceeds the scale of Wilddash2 annotations (5509 vs. 5000). Moreover, in contrast both to Wilddash2 and to other recent dense semantic perception datasets with adverse conditions [62, 63], ACDC features cross-condition image-level correspondences with normal-condition reference images as well as a specialized annotation protocol which hinges exactly on the aforementioned correspondences to allow the assignment of legitimate, reliable semantic labels to indiscernible image regions which would otherwise be impossible to label.
@@
Instance segmentation and panoptic segmentation distinguish instances in the images compared to the semantic segmentation task. Instance segmentation aims to assign a segmentation mask for each object of interest in the image, while the panoptic segmentation task encompasses masks for both stuff and things classes as the combination of semantic segmentation and instance segmentation. Both of these tasks are applied in various real-world applications such as robotics, healthcare and geoscience. For instance segmentation, many works either follow the detect-then-segment pipeline [87] or explore the generation of a dynamic number of instance masks directly with clustering or dynamic kernels. The pioneering work of Mask R-CNN [87] generates instance masks based on object bounding box proposals. The following specialized instance segmentation works such as Cascaded Mask R-CNN, HTC and Detectors mainly focus on extracting better feature representations and generating more accurate object proposals. For panoptic segmentation, early works strive to combine semantic segmentation models and instance segmentation models to predict unified panoptic masks. Recent works begin to view the panoptic segmentation task in a unified perspective and formulate the stuff and things segmentation as a set prediction problem. The current state of the art for panoptic segmentation includes i.a. PanopticFPN and Panoptic-Deeplab with specialized instance prediction branches and semantic prediction branches, K-Net with dynamic kernels, and MaskFormer and Mask2Former with transformer architectures. Although current instance segmentation and panoptic segmentation models obtain impressive improvement on the popular Cityscapes benchmark, we show that these models generalize poorly to ACDC. The adverse images in ACDC pose a new challenge for instance and panoptic segmentation models and encourage further progress in the field.
@@
Adaptation of object detection networks from a labeled source domain to another unlabeled target domain is also an active research field. Currently, there are mainly two categories of UDA for object detection methods: domain alignment and self-training. Domain alignment strives to bridge the domain gap by minimizing the domain discrepancy through style transfer, adversarial training and graph matching. Self-training relies on pseudo-labels to extract rich knowledge contained in the target domain and presents promising performance. Although all aforementioned strategies have been proven effective in alleviating the damage of domain shift, due to the lack of datasets with rich adverse conditions, these methods are mainly evaluated on synthetic-to-real, cross-city and cross-camera settings. The normal-to-adverse adaptation, which is of high value for intelligent driving systems, is rarely examined. ACDC provides a large-scale dataset for adverse conditions and promotes the progress of unsupervised and weakly supervised object detection methods under adverse conditions.
@@ 3 ACDC Dataset
We base the design of ACDC on the same general principles as seminal normal-condition datasets [1] and adapt the collection and annotation process to fit better the adverse condition setting at hand. An overview of this process is shown in Fig. 2.
@@ 3.1 Collection
Thus, we recorded several days of video in Switzerland by driving around in a car, primarily in urban areas but also on highways and in rural regions. In order to have a clear domain separation between different adverse conditions, we use the following criterion for the adverse-condition recordings: each recording takes place under only one type of adversity from a set of four types, i.e., fog, nighttime, rain, and snow. For example, our foggy recordings are performed at daytime and without rain or snow. For snow, both snowfall and snow cover on the ground are admissible. Moreover, we keep for further processing only the parts of the adverse-condition recordings that correspond to an intense presence of the respective condition, so as to maximize the domain shift from normal conditions as well as domain adversity.
@@ 3.2 Correspondence Establishment
Our camera also provides GPS readings, which allow us to establish image-level correspondences between adverse-condition and normal-condition recordings. In particular, for each adverse-condition recording, we perform a normal-condition recording along exactly the same route. We then use the sequences of GPS measurements of the two recordings to perform a global dynamic-programming-based matching of the adverse GPS sequence to the normal one, where the objective is defined by the Euclidean distances of matched pairs of GPS samples. Our global matching handles routes with loops better than simple nearest neighbors. Each adverse-condition frame is then matched to a normal-condition frame based on the corresponding matched samples of the GPS sequences, as shown in the left part of Fig. 2.
@@ 3.4 Semantic Annotation
Our annotation protocol for the 4006 adverse-condition images consists of two cascaded annotation stages. At stage 1, a semantic labeling draft is manually produced from the adverse-condition image , in which pixels that cannot be unquestionably assigned to a single semantic class are left unlabeled. At stage 2, the corresponding normal-condition image and the adverse-condition video from which was extracted are used to augment and finalize the annotation. In particular, the annotator can assign a legitimate label to pixels that were left unlabeled in stage 1 and correct pixels that were incorrectly labeled in stage 1. Annotation during stage 2 using the corresponding normal-condition image and the neighboring frames of the adverse-condition video does not assume or require an exact match of the 6D camera pose between the annotated image and the auxiliary images or frames. In particular, the transfer of labels from the auxiliary images to the image under annotation is done through manual inspection by the annotators rather than automatically. We have explicitly instructed the annotators to be highly conservative in this manual transfer given potential mismatch of pose and dynamic image content, which is why 7.5% of the pixels in the annotated images are left unlabeled even after stage 2. More specifically, pixels that remain unclear in stage 2 are left unlabeled and are not used for training or evaluation.
@@
The 4006 fine pixel-level annotations of ACDC were created by a professional team of annotators to ensure high-quality ground truth. Annotators were asked to be conservative in labeling pixels in both stages, so as to minimize errors. Both the initial draft from stage 1 and the final annotation from stage 2 passed through quality control (QC). QC was performed via manual inspection of each annotation by a second person (B), different from the one who created it (A). A and B were given incentives to minimize errors and maximize identified errors, respectively. A third independent person (C) verified whether each error identified by B was factual. The total average time required for annotating a single adverse-condition image was 3.3 h, with 25% of it spent on QC. The semantic annotation of the 1503 normal-condition images is conducted in the standard way by using only the input normal-condition image.
The class specifications of ACDC are directly inherited from Cityscapes. In particular, we annotate the 19 evaluation classes of Cityscapes, which include the most common and traffic-related objects in driving scenes. Objects that belong to classes outside this set receive a fall-back label and are not used for training or evaluation. This choice of classes provides full compatibility of ACDC to Cityscapes and other normal-condition datasets for semantic segmentation [2, 5]. Detailed annotation statistics are presented in Fig. 1. An example of our two-stage annotation protocol is shown in Fig. 2 for a snowy image. Note the assignment of a region in the lower right part of the image that is unlabeled at stage 1 (top right) to the road label at stage 2 (middle right), thanks to the clear view from the normal-condition image.